• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 01:53:40 on September28,2005

    Entry number 156299

    keyword=more Q2 scanner studies, part I

    Part I of Q2 scanner runs with various magnet settings.

    Run 5425 -- everything normal (to get a baseline), i.e. sweeper is
    on, beam is 36 uA. After about 13K events we see R-sept at 7.69K,
    so we ask for 30 uA, then ask for beam off. Stop run. Too hot !

    BTW, now I think I know why the voice alarms didn't work this
    morning when we tripped: someone had set the alarm level to
    "greater than 7.695", which I think is too high...

    Run 5426 -- sweeper still on but we reduce current to 25 uA since
    it will tend to overheat septum when we do subsequent tests.

    01:22 sweeper turned off

    Run 5427 -- scan with sweeper off
    Again the septum gets too hot ! 7.72K. So we turn beam off,
    stop run 5427 at event 15K. After a few minutes cooling we
    ask for 10 uA.

    Run 5428 -- sweeper still off. 10 uA now.
    Fig 1 shows the scan results for run 5428. Shoulder qualitatively
    the same as before.

    Run 5429 -- sweeper back on. 111Amps. Still 10uA beam. We wanted
    to repeat the baseline with this lower beam current.

    More to come ...


    FIGURE 1